Express the following as the sum of two odd primes.
(a) 44
(b) 36
(c) 24
(d) 18

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) 44 = 13 + 31 (b) 36 = 17 + 19 (c) 24 = 7 + 17 (d) 18 = 7 + 11